# Presenting Interval Pomsets with Interfaces

### v1

@@ 1 Introduction
Another class of pomsets which are important in concurrency theory are interval orders [14, 15]. These are pomsets whose events may be represented as intervals on the real line and have found their place in relativity [29], concurrency theory [22, 20, 21], and distributed computing [24, 25, 5]. Their algebraic theory, however, is less developed. Starting in [9], a notion emerged that for the first, ipomsets in their full generality may not be needed for concurrency but interval orders suffice, and secondly, that ipomsets might provide a suitable algebraic theory for interval orders. Picking up on ideas in [12, 2, 1] and based on the antichain representations of [21], the purpose of this paper is to develop such an algebraic theory of interval orders.
@@ 2.1 Pomsets with interfaces
We may add subscripts “${}_{P}$” to the elements above if necessary and omit any empty substructures from the signature. We will also often use the notation ${}_{S}P_{T}$ instead of $(P,{<},{\dashrightarrow},S,T,\lambda)$ if no confusion may arise.
{tikzpicture}
@@
Figure 1 depicts an ipomset $P=\{x_{1},x_{2},x_{3},x_{4}\}$ with four events labelled $\lambda(x_{1})=a$, $\lambda(x_{2})=b$, $\lambda(x_{3})=c$, and $\lambda(x_{4})=d$. (We do not show the identity of events, only their labels.) Its precedence order is given by $x_{1}<x_{2}$, $x_{3}<x_{2}$ and $x_{3}<x_{4}$ and its event order by $x_{1}\dashrightarrow x_{3}$, $x_{1}\dashrightarrow x_{4}$ and $x_{2}\dashrightarrow x_{4}$. The sources are $S=\{x_{3}\}$ and the targets $T=\emptyset$. (We denote these by “ $\bullet$ ”.) We think of events in $S$ as being already active at the beginning of $P$, and the ones in $T$ (here there are none) continue beyond the ipomset $P$.
@@
The source interface of $P$ is the conclist $S_{P}=(S,{\dashrightarrow_{\upharpoonleft S\times S}},\lambda_{\upharpoonleft S})$ where “${}_{\upharpoonleft}$” denotes restriction; the target interface of $P$ is the conclist $T_{P}=(T,{\dashrightarrow_{\upharpoonleft T\times T}},\lambda_{\upharpoonleft T})$. We call a starter or terminator elementary if $|S|=1$, resp. $|T|=1$, that is if it starts or terminates exactly one event.
@@
  • by St the set of starters ${}_{S}U_{U}=(U,{\dashrightarrow},S,U,\lambda)$;
  • by Te the set of terminators ${}_{U}U_{T}=(U,{\dashrightarrow},U,T,\lambda)$;
@@
    <=({${<}=\big{(}\{(i(x),i(y)\mid x<_{P}y\}\cup\{j(x),j(y)\mid x<_{Q}y\}\cup\{(i(x),j(y)\mid x\in P\setminus T_{P},y\in Q\setminus S_{Q}\}\big{)}^{+}$, where $i:P\to R$ and $j:Q\to R$ are the injections and ${}^{+}$ denotes transitive closure;
@@
  • identities are $\textup{{id}}_{U}={}_{U}U_{U}\in\text{{{iiPoms}}}_{\simeq}(U,U)$. ∎
@@ 2.2 Starters and terminators
such that $T_{P_{i}}=S_{P_{i+1}}$ for all $i=1,\dots,n-1$. Such words are called coherent in [2]. Note that $P_{1}\dots P_{n}$ is coherent iff the gluing $P_{1}*\dots*P_{n}$ is defined.
@@
Let $\sim$ be the congruence on $\textup{{Coh}}_{\simeq}$ generated by the relationsZ
@@
$\textup{{id}}_{U}$ $\sim{}_{U}U_{U}\qquad$ $(U\in\square).$
@@ 2.3 Step sequences
$P=\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\\ \vbox{\hbox{\tiny$\bullet$}}d\end{matrix}\right]:$
@@
it first starts $a$, then terminates $c$, then starts $d$ and terminates $a$, then starts $b$ and finally terminates both $b$ and $d$. The dense step decompositions of $P$ are
@@
$P$ =[$a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}d\end{matrix}\right]$
@@
=[$a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}d\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}d\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\end{matrix}\right],$
which differ only in the order in which $b$ and $d$ are terminated at the end.
@@
The category $\text{{{iiPoms}}}_{\simeq}$ is freely generated by the directed multigraph Ω¯subscriptnormal-¯normal-$\bar{\Omega}_{\simeq}$ using gluing composition.
@@ 3 Subsumptions in Step Sequences
In this section we extend the equivalence between ipomsets and step sequences from Thm. 2.1 to also cover subsumptions.
@@
$\preceq$ is a partial order on $O$ and has a minimal element $(x_{0},y_{0})$. Let $R$ be defined as $P$, but with ${<_{R}}={<_{P}}\setminus\{(x_{0},y_{0})\}$ and ${\dashrightarrow_{R}}={\dashrightarrow_{Q}}$, then $\textup{{id}}:P\to R$ is an elementary subsumption and $f:R\to Q$ a subsumption.
@@
Reflexivity, anti-symmetry and transitivity of $\preceq$ come from $\leq_{P}$. As $O$ is finite, there exists a minimal element.
@@
Irreflexivity and anti-symmetry are given by $<_{P}$. For transitivity, let $x<_{R}y$ and $y<_{R}z$. Then $x<_{P}z$ and we have to show that either $x\neq x_{0}$ or $z\neq y_{0}$. Assuming $(x,z)=(x_{0},y_{0})$, there would be $y$ such that $x_{0}<_{R}y<_{R}y_{0}$ which contradicts the minimality of $(x_{0},y_{0})$ in the order $\preceq$.
@@
Let us now introduce some useful notation for starters and terminators to more clearly specify the conclists of events which are started or terminated. For $U\in\square$ and $A,B\subseteq U$ we write ${{}_{A\!}}{{\uparrow}U}{}={}_{U\setminus A}U_{U}$ for the starter which starts the events in $A$ and ${U}{\downarrow}_{B}={}_{U}U_{U\setminus B}$ for the terminator which terminates the events in $B$. In the following definition we express the result of transpositions of elements of a coherent word $w=P_{1}\dots P_{n}\in\textup{{Coh}}_{\simeq}$.
@@
{tikzpicture}
Figure 2: Interval representations of several ipomsets, cf. Ex. 3.
@@
Let $f:P\to Q$ be a subsumption and $P=P_{1}*\dots*P_{n}$, $Q=Q_{1}*\dots*Q_{n}$ dense step decompositions. There exist dense words $u_{1},\dots,u_{m}\in\textup{{Coh}}_{\simeq}$ for some $m\geq 0$, and indices $i_{1},\dots,i_{m-1}\in\{1,\dots,n-1\}$ such that $u_{1}=w_{1}=P_{1}\dots P_{n}$, $u_{m}=w_{2}=Q_{1}\dots Q_{n}$ and, for all $k$, $u_{k+1}=\tau_{i_{k}}(u_{k})$.
@@
Let $<_{e}$ be the relation on $\textup{{Coh}}_{\simeq}$ defined by $w_{1}<_{e}w_{2}$ if there is an index $i$ such that $w_{2}=\tau_{i}(w_{1})$ and $P_{i}P_{i+1}={{}_{A\!}}{{\uparrow}U}{}{U^{\prime}}{\downarrow}_{B}$ with $A\cap B=\emptyset$ (third case of Def. 3). Denote by the same symbol ${<}_{e}$ the relation induced in the quotient $\textup{{Coh}}_{\sim}$ and by ${\leq}={<_{e}^{*}}$ the reflexive, transitive closures.

### v2

@@ 1 Introduction
Another class of pomsets which are important in concurrency theory are interval orders [15, 16]. These are pomsets whose events may be represented as intervals on the real line and have found their place in relativity [31], concurrency theory [23, 21, 22], and distributed computing [25, 26, 6]. Their algebraic theory, however, is less developed. Starting in [10], a notion emerged that for the first, ipomsets in their full generality may not be needed for concurrency but interval orders suffice, and secondly, that ipomsets might provide a suitable algebraic theory for interval orders. Picking up on ideas in [13, 2, 3, 1] and based on the antichain representations of [22], the purpose of this paper is to develop such an algebraic theory of interval orders.
@@ 2.1 Pomsets with interfaces
The precedence order is the ”usual” order: $a<b$ means that $a$ ended before the beginning of $b$. The event order is necessary in order to define gluing composition, see below.
@@
We may add subscripts “P” to the elements above if necessary and omit any empty substructures from the signature. We will also often use the notation ${\vphantom{P}}_{S}P_{T}$ instead of $(P,{<},{\dashrightarrow},S,T,\lambda)$ if no confusion may arise.
$a\vphantom{d}$$c\vphantom{d}$bullet $b\vphantom{d}$$a\vphantom{d}$
@@
Figure 1 depicts an ipomset $P=\{x_{1},x_{2},x_{3},x_{4}\}$ with four events labelled $\lambda(x_{1})=\lambda(x_{4})=a$, $\lambda(x_{2})=b$ and $\lambda(x_{3})=c$. (We do not show the identity of events, only their labels.) Its precedence order is given by $x_{1}<x_{2}$, $x_{3}<x_{2}$ and $x_{3}<x_{4}$ and its event order by $x_{1}\dashrightarrow x_{3}$, $x_{1}\dashrightarrow x_{4}$ and $x_{2}\dashrightarrow x_{4}$. The sources are $S=\{x_{3}\}$ and the targets $T=\emptyset$. (We denote these by “ $\bullet$ ”.) We think of events in $S$ as being already active at the beginning of $P$, and the ones in $T$ (here there are none) continue beyond the ipomset $P$.
@@
A conclist is, thus, a pomset of the form $(P,\emptyset,\dashrightarrow,\emptyset,\emptyset,\lambda)$; to ease notation, we will omit the $\emptyset$ from the notation.
The source interface of an ipomset $(P,{<},{\dashrightarrow},S,T,\lambda)$ as above is the conclist $S_{P}=(S,{\dashrightarrow_{\upharpoonleft S\times S}},\lambda_{\upharpoonleft S})$ where “” denotes restriction; the target interface of $P$ is the conclist $T_{P}=(T,{\dashrightarrow_{\upharpoonleft T\times T}},\lambda_{\upharpoonleft T})$.
We call a starter or terminator elementary if $|S|=1$, resp. $|T|=1$, that is if it starts or terminates exactly one event. In the following, a discrete ipomset will be represented by a vertical ordering of its elements following the event order, with elements in the source (resp. target) set preceded (resp. followed) by the symbol $\bullet$ .
Example 2
The ipomset ($(\{x_{1},x_{2},x_{3}\},\emptyset,x_{1}\dashrightarrow x_{2}\dashrightarrow x_{3},\{x_{1},x_{2}\},\{x_{2},x_{3}\},$$\{(x_{1},a),(x_{2},b),(x_{3},c)\})$ can be represented by [$a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}b\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}c\vbox{\hbox{\tiny$\bullet$}}\end{smallmatrix}\right]$.
@@
  • by St the set of starters ${\vphantom{U}}_{S}U_{U}=(U,{\dashrightarrow},S,U,\lambda)$;
  • by Te the set of terminators ${\vphantom{U}}_{U}U_{T}=(U,{\dashrightarrow},U,T,\lambda)$;
@@
    <=({${<}=\big{(}\{(i(x),i(y))\mid x<_{P}y\}\cup\{(j(x),j(y))\mid x<_{Q}y\}\cup\{(i(x),j(y))\mid x\in P\setminus T_{P},y\in Q\setminus S_{Q}\}\big{)}^{+}$, where $i:P\to R$ and $j:Q\to R$ are the injections and + denotes transitive closure;
@@
  • identities are $\textup{{id}}_{U}={\vphantom{U}}_{U}U_{U}\in\text{{{iiPoms}}}_{\simeq}(U,U)$. ∎
@@ 2.2 Starters and terminators
such that $T_{P_{i}}=S_{P_{i+1}}$ for all $i=1,\dots,n-1$. Such words are called coherent in [3, 2]. Note that $P_{1}\dots P_{n}$ is coherent iff the gluing $P_{1}*\dots*P_{n}$ is defined.
@@
Let $\sim$ be the congruence on $\textup{{Coh}}_{\simeq}$ generated by the relations
@@
$\textup{{id}}_{U}$ $\sim{\vphantom{U}}_{U}U_{U}\qquad$ $(U\in\square).$
@@ 2.3 Step sequences
$P=\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\\ \vbox{\hbox{\tiny$\bullet$}}a\end{matrix}\right]:$
@@
it first starts the first $a$, then terminates $c$, then starts the second $a$, terminates $a$, then starts $b$ and finally terminates both $b$ and the second $a$. The dense step decompositions of $P$ are
@@
$P$ =[$a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}d\end{matrix}\right]$
@@
=[$a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}c\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\\ \phantom{\vbox{\hbox{\tiny$\bullet$}}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\phantom{\vbox{\hbox{\tiny$\bullet$}}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}a\vbox{\hbox{\tiny$\bullet$}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\vbox{\hbox{\tiny$\bullet$}}\\ \vbox{\hbox{\tiny$\bullet$}}a\phantom{\vbox{\hbox{\tiny$\bullet$}}}\end{matrix}\right]\left[\begin{matrix}\vbox{\hbox{\tiny$\bullet$}}b\end{matrix}\right],$
which differ only in the order in which $b$ and $a$ are terminated at the end.
@@
The category $\text{{{iiPoms}}}_{\simeq}$ is generated by the directed multigraph $\bar{\Omega}_{\simeq}$ using gluing composition under the identities (2).
@@ 3 Subsumptions in Step Sequences
In this section, we extend the equivalence between ipomsets and step sequences from Thm. 2.1 to also cover subsumptions.
@@
$\preceq$ is a partial order on $O$ and has a maximal element $(x_{0},y_{0})$. Let $R$ be defined as $P$, but with ${<_{R}}={<_{P}}\setminus\{(x_{0},y_{0})\}$ and ${\dashrightarrow_{R}}={\dashrightarrow_{Q}}$, then $\textup{{id}}:P\to R$ is an elementary subsumption and $f:R\to Q$ a subsumption.
@@
Reflexivity, anti-symmetry and transitivity of $\preceq$ come from $\leq_{P}$. As $O$ is finite, there exists a maximal element.
@@
Irreflexivity and anti-symmetry are given by $<_{P}$. For transitivity, let $x<_{R}y$ and $y<_{R}z$. Then $x<_{P}z$ and we have to show that either $x\neq x_{0}$ or $z\neq y_{0}$. Assuming $(x,z)=(x_{0},y_{0})$, there would be $y$ such that $x_{0}<_{P}y<_{P}y_{0}$ which contradicts the maximality of $(x_{0},y_{0})$ in the order $\preceq$.
@@
Let us now introduce some useful notation for starters and terminators to more clearly specify the conclists of events which are started or terminated. For $U\in\square$ and $A,B\subseteq U$ we write ${{}_{A\!}}{{\uparrow}U}{}={\vphantom{U}}_{U\setminus A}U_{U}$ for the starter which starts the events in $A$ and ${U}{\downarrow}_{B}={\vphantom{U}}_{U}U_{U\setminus B}$ for the terminator which terminates the events in $B$. In the following definition we express the result of transpositions of elements of a coherent word $w=P_{1}\dots P_{n}\in\textup{{Coh}}_{\simeq}$.
@@
[…]
Figure 2: Interval representations of several ipomsets, cf. Ex. 4.
@@
Let $f:P\to Q$ be a subsumption and $P=P_{1}*\dots*P_{n}$, $Q=Q_{1}*\dots*Q_{n}$ dense step decompositions. Then there exist dense words $u_{1},\dots,u_{m}\in\textup{{Coh}}_{\simeq}$ for some $m\geq 0$, and indices $i_{1},\dots,i_{m-1}\in\{1,\dots,n-1\}$ such that $u_{1}=w_{1}=P_{1}\dots P_{n}$, $u_{m}=w_{2}=Q_{1}\dots Q_{n}$ and, for all $k$, $u_{k+1}=\tau_{i_{k}}(u_{k})$.
@@
Let $<_{e}$ be the relation on $\textup{{Coh}}_{\simeq}$ defined by $w_{2}<_{e}w_{1}$ if there is an index $i$ such that $w_{2}=\tau_{i}(w_{1})$ and $P_{i}P_{i+1}={{}_{A\!}}{{\uparrow}U}{}{U^{\prime}}{\downarrow}_{B}$ with $A\cap B=\emptyset$ (third case of Def. 3). Denote by the same symbol ${<}_{e}$ the relation induced in the quotient $\textup{{Coh}}_{\sim}$ and by ${\leq}={<_{e}^{*}}$ the reflexive, transitive closures.